OpenStudy (anonymous):

Mr Castle Shares out 64 marbles Emily and Asif in the ratio 7:1How many Marbles does each child get? Emily gets...... Marbles and Asif Gets.....Marbles

OpenStudy (shubhamsrg):

let emily have E marbles, asif have A marbles given that E/A=7/1 and E+A = 64 2 eqns, 2 variables, you know what to do,best of luck! :P